Zany Brainy Corp. manufactures two types of lamps, Home size and Industry size lamps. Company produces 15,000 Home size lamp and 5,000 Industry size lamps. Following is the data provided about these two products:

Home size lamp

Industry size lamp

Total

Direct Materials per unit

$7

$12

Direct manufacturing labour rate per unit

$12.00

$12.00

Direct labour hours per unit

0.50

0.6

Set up costs

$120,000

Lamps produced per patch

200

50

Set up hour per patch

1

3

Quality control costs

$160,000

Total indirect manufacturing costs

$280,000

Annual production

15,000

5,000

Page Break

  1. The company’s simple (traditional) costing system uses a single indirect-cost pool and allocates costs to the two lamps on the basis of direct manufacturing labour-hours. Under this tradition system, determine total product cost per unit of each of the company's two products. (10 marks)

  2. The company’s ABC system uses two indirect-cost pools. One cost pool is for set up costs and the other for quality control. Zany Brainy allocates set up costs to the two lamps based on setup labour-hours and quality control costs based on direct manufacturing labour-hours. Under ABC system, determine total product cost per unit of each of the company's two products. (25 marks)
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Zany Brainy Corp.
Workings for Answer 1 Home Size Industry Size Total Note
Direct Labor per unit                           0.50                   0.60 A
Units produced                 15,000.00            5,000.00 B
Total Direct Labor Hours                   7,500.00           3,000.00        10,500.00 C=A*B
Total overhead      280,000.00 D
Plant wide OH Rate                26.67 E=D/C
Allocated OH              200,000.00         80,000.00      280,000.00 F=C*E
Allocated OH per unit                        13.33                 16.00 G=F/B
Answer 1 Home Size Industry Size Note
Direct Materials                           7.00                 12.00
Direct Labor                        12.00                 12.00
Allocated OH per unit                        13.33                 16.00 See G
Unit product cost- Traditional                        32.33                 40.00
Workings for Answer 2 Home Size Industry Size Total Note
Units produced                 15,000.00            5,000.00 See B
Lamps produced per patch                      200.00                 50.00 H
Number of patches                        75.00               100.00 I=B/H
Set up hours per patch                           1.00                   3.00 J
Total Set up hours                        75.00               300.00              375.00 K=I*J
Calculation of ABC rates for overhead L M N=L/M
Activity cost pool Activity Measure Cost Activity usage ABC rates
Setup costs Setup hours       120,000.00              375.00             320.00
Quality Control costs Direct Labor Hours       160,000.00        10,500.00               15.24
Total Cost assigned       280,000.00
Home Size See L O P=L*O
Activity cost pool Activity Measure ABC Rates Activity used Cost assigned
Setup costs Setup hours               320.00                75.00        24,000.00
Quality Control costs Direct Labor Hours                 15.24           7,500.00     114,285.71
Total Cost assigned     138,285.71 Q
Units produced        15,000.00 See B
Costs per unit                  9.22 R=Q/B
Industry Size See L S T=L*S
Activity cost pool Activity Measure ABC Rates Activity used Cost assigned
Setup costs Setup hours               320.00              300.00        96,000.00
Quality Control costs Direct Labor Hours                 15.24           3,000.00        45,714.29
Total Cost assigned     141,714.29 U
Units produced          5,000.00 See B
Costs per unit               28.34 V=U*B